OpenStudy (anonymous):

Write the following complex number in polar form. Express all angles in degrees rounded to the nearest tenth. –3

zepdrix (zepdrix):

\[\Large\rm -3=\color{royalblue}{3}\cdot\color{orangered}{(-1)}=\color{royalblue}{r}\cdot\color{orangered}{e^{i?}}\]

zepdrix (zepdrix):

You have learned about the exponential form, or no?

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{3}\cdot\color{orangered}{(-1)}=\color{royalblue}{r}\color{orangered}{(\cos \theta+i \sin \theta)}\]Maybe this breakdown is unnecessary. -3 is along the real axis. So what angle is that? :)

zepdrix (zepdrix):